Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 01

1. Section 8.7: Do as many as you can the odd number problems in problem 35-65.

2. Section 8.7: Evaluate

∫_{0}^{1} \frac{1}{x^{p}} dx = lim_{ε→0} ∫_{ε}^{1} \frac{1}{x^{p}} dx

for 0 < p < 1, p = 1 and p > 1 respectively.

3. Section 10.3: Problems 27, 31, 33, 37, 53, 55.